OpenStudy (anonymous):

The number 0 can be classified as A. an integer. B. a real number. C. a whole number. D. all of the above

OpenStudy (anonymous):

All rational numbers are also A. real numbers. B. natural numbers. C. integers. D. whole numbers.

OpenStudy (anonymous):

0.5 is what type of number? A. an integer B. a natural number C. a rational number D. a whole number

OpenStudy (anonymous):

for the first one: I say all of the above because 0 is a neutral number

OpenStudy (anonymous):

All integers are also A. rational numbers. B. whole numbers. C. natural numbers. D. irrational numbers.

OpenStudy (anonymous):

Which of these number is an integer? A. −4.0 B. 1.50 C. D.

OpenStudy (anonymous):

A. for the last one

OpenStudy (anonymous):

the rest of them I don't know

OpenStudy (anonymous):

What type of number is ? 163 A. an integer B. a rational number C. a whole number D. an irrational number

OpenStudy (anonymous):

either a rational number or a whole number for the number 163

OpenStudy (anonymous):

can you medal me?

OpenStudy (anonymous):

can u answer all my q

OpenStudy (anonymous):

I'm sorry I don't know the rest

OpenStudy (anonymous):

all integers are also irrational

OpenStudy (error1603):

k

OpenStudy (anonymous):

I do know that one

OpenStudy (anonymous):

0.5 is a natural number

OpenStudy (anonymous):

All rational numbers are also whole numbers

OpenStudy (anonymous):

there I think that is all of them

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

did you medal me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k thx

OpenStudy (error1603):

k

OpenStudy (xmissalycatx):

Which questions do you still have? Just type them below and I'll help you.

OpenStudy (anonymous):

Marcos drove twice as far as Candice. Together, they drove 66 miles. Which equation can be used to find the number of miles Candice drove? A. 3x = 66 B. C. 3x = 66 + x D.

OpenStudy (anonymous):

The length of a swimming pool is 22 ft. The width is 40 ft. If the volume of the pool is 2,500 ft³, which equation below could be used to find the depth of the pool? A. (22 + 40)x = 2500 B. 22 + 40 + x = 2500 C. 66x = 2500 D. 22 40 x = 2500

OpenStudy (anonymous):

A secretary at the police department makes $8 less per hour than the dispatcher. The combined hourly wage of the secretary and the dispatcher is $40. Which equation could be used to find the hourly wage of the dispatcher? A. 2x + 8 = 40 B. 2x − 8 = 40 C. 8 − 2x = 40 D. (8 − 2)x = 40

OpenStudy (anonymous):

A 180 ft long board is cut into 3 pieces. The second piece of board is twice as long as the first. The third piece of board is six times as long as the first. What is the length of the first piece of the board? A. 20 ft B. 40 ft C. 60 ft D. 120 ft

OpenStudy (xmissalycatx):

For the first one you sent about driving.. what are the other two answer choices?

OpenStudy (anonymous):

3x=-66

OpenStudy (xmissalycatx):

The second one is D because you use the formula V= l(w)(h)

OpenStudy (anonymous):

ok

OpenStudy (xmissalycatx):

Third one is D as well.

OpenStudy (anonymous):

wich one the sectary

OpenStudy (xmissalycatx):

The last one is A

OpenStudy (xmissalycatx):

The one with Marcos and Candice, I don't know because there's 2 answer choices missing.. But the other 3 in order are D D A

OpenStudy (anonymous):

A secretary at the police department makes $8 less per hour than the dispatcher. The combined hourly wage of the secretary and the dispatcher is $40. Which equation could be used to find the hourly wage of the dispatcher? A. 2x + 8 = 40 B. 2x − 8 = 40 C. 8 − 2x = 40 D. (8 − 2)x = 40

OpenStudy (anonymous):

and this one

OpenStudy (xmissalycatx):

That one was the first one I told you. It is D.